NMSU 1 meter telescope: design and performance

Jon Holtzman, NMSU July 10, 2013

Complete document in a single PostScript file.

1 Telescope: optical/mechanical design

• Telescope is an alt-az Ritchey-Chretien design and produces an F/6.06 beam.

• Basic design optical parameters (units are inches)

The telescope mirrors were refigured in 2002 because we strongly suspected that the surfaces were not particularly good (which was confirmed by optical testing before repolishing). During repolishing the followin optical prescription was adopted:

NAME  RADIUS THICKNESS DIAMETER CONIC
Primary  -200  -62.1  40  -1.202461
Secondary -128.8185  47.50  -8.38632
Tertiary  Inf  44.784

Focal plane

The following discussion describes how this prescription was arrived at.

• For the original optical design, our best idea of the final prescription is taken from a document provided by Kurt Anderson (values labelled “Preferred” for 20 Dec 1993):

NAME  RADIUS THICKNESS DIAMETER CONIC
Primary  -200  -62.648  40  -1.202461
Secondary -127.059  49.00  -8.434993
Tertiary  Inf  41.648

Focal plane

We have essentially no information as to how the as-built optical parameters match these design parameters.

When we built our new guide box, our best measured distance from tertiary to rotator mount was 30 inches (however, see below!). Our old guide box/filter wheel had a total thickness of 9.875 inches, for a total tertiary to focus distance of 39.875 inches.

Our new guide box/filter wheel was designed to increase the tertiary to focus distance by a couple of inches to better match the design prescription (under the assumption that the primary-secondary separation matched the design value, see below). The guide box has a thickness of 9.25 inches, the filter wheel, 1.313 inches, and our best measurement for the optical distance of the Princeton CCD from the front mounting surface is 1.125 inches, for a total tertiary to focus distance of 41.688 inches.

The prescription I was then using for ray tracing in Zemax is:

NAME  RADIUS THICKNESS DIAMETER CONIC
Primary  -200  -62.66  40  -1.202461
Secondary  -127.059  48.6585  -8.434993
Tertiary  Inf  30
Rotator mount  11.93

Focal plane

where the focal plane location and primary-secondary separation were allowed to float to get best image quality.

We later switched to using an Apogee AP7P camera, which has the CCD 1.10 inches behind the front mounting surface (as measured by microscope); we added an adapter plate of width .375 inches (for a reason that is not totally recalled!).

• Measured separations. Several (re)measurements of separations were made in 2002 in preparation for mirror repolishing.

Several of these measurements were made from the edge of the primary to the center of another element (tertiary or secondary). To correct for the distance to the primary vertex requires a correction from the hypotenuse of a triangle to the on-axis distance, plus a correction for the vertical distance between the primary vertex and the edge of the mirror. For the hypotenuse correction, a diameter of the front surface of the mirror was measured (approximately) to be 41 inches; note that using a smaller diameter leads to some inconsistencies in measurements (see JCB 6/02 and JH 6/02). The vertex “sag” was computed using a radius of curvature of 200 inches, resulting in a 1 inch sag.

JCB 4/02: measured distance from primary edge to secondary center as 166.2+/-0.3 cm (65.43 inches); correcting for hypotenuse gives (62.14+/- 0.12 inches). Distance from vertex to ”edge center” adds another inch based on radius of curvature, to give 63.14. Using geometric center of tertiary, measured secondary-tertiary to be 1227 cm (48.31 inch). Tertiary to rotator mount measured outside of bearing to be 79.4cm (31.25”), rotator mount to filter mount 27.3 cm (10.75”)

JH/DS 5/20: measure primary-secondary with tertiary removed. Method 1: mirror to mirror cover lip 2”, lip to tertiary mount 5”, tertiary mount to secondary center 55.5”, total 62.5 + distance from primary vertex to spot on primary out at edge of central hole. Method 2: secondary center to primary edge (as JCB) gives 65”, correcting for hypotenuse gives 61.68; add another inch as above to get 62.68. Using plumb line to determine tertiary center, get 47.56 ” secondary-tertiary. Tertiary to CCD mount measured through the hole to give 42.185”

JCB 6/02: distance from primary edge to tertiary center 24.75 inches, corrected for hypotenuse gives 13.87, plus 1 for sag gives 14.87 inches. Distance from tertiary to secondary is 47.71. Distance from primary edge to secondary center is 64.81, corrected for hypotenuse is 61.48, plus 1 for sag is 62.48. (Note for consistency check: 14.87 + 47.71 = 62.58) Tertiary to CCD mount is 42.19 inches. CCD mounting plate measured to be .375 inches. Apogee web site gives 1.161 inches for distance from front of camera to CCD.

JH 6/02: distance from primary edge to tertiary center 24.625 inches, corrected for hypotenuse gives 13.64, plus 1 for sag gives 14.64 inches. Distance from tertiary to secondary is 47.875. Distance from primary edge to secondary center is 64.875, corrected for hypotenuse is 61.55, plus 1 for sag is 62.55. (Note for consistency check: 14.64 + 47.875 = 62.515)

JH/JCB 6/02: microscope gives distance from front window of Apogee to CCD surface is 0.802 inches. Micrometer give 0.682 inches from front mounting surface to front window. Mounting plate thickness 0.375 inches.

Summary (numbers in parentheses are derived from other measurements) :

distance  design  Zemax JCB 4/02 JH/DS 5/02  JCB 6/02 JH 6/02
primary-secondary  62.648 62.66 63.14  62.5+,62.68  62.48  62.55
secondary-tertiary  49  48.66 48.31  47.56  47.71  47.875
primary-tertiary  (13.648) (14)  (14.83)  (15.12)  14.87  14.64
tertiary-rotator mount  31.25  (31.622)  (31.627) (31.622)
rotator mount-filter mount  10.75  10.563(design) 10.563  10.563
tertiary-filter mount  (42)  42.185  42.19  42.185
CCD mounting plate thickness  0.375  .375  .375
front of camera to CCD  1.161  1.109  1.109
tertiary-focal plane  41.648 41.93 (43.54)  (43.72)  (43.67)  (43.67)
(secondary-focal plane)  90.648 90.59 (91.84)  (91.28)  (91.38)  (91.545)

Adopt as-built:

primary-tertiary  14.8
tertiary-secondary  47.875
tertiary-filter mount 42.185
mounting plate  .375
camera surface-CCD 1.11

In measured configuration, secondary has range to move approximately 0.125 inch away from primary, and 1 inch towards primary. Primary support could be modified to move primary towards secondary (e.g. increase pad thickness), but not necessarily further away.

For new design, modify distances conservatively given freedom to adjust primary towards secondary, secondary towards tertiary, and CCD away from tertiary:

primary-tertiary  14.6
tertiary-secondary  47.5
tertiary-filter mount 42.2
mounting plate  .375
camera surface-CCD 1.11

• sizes of optics. JCB 4/02 measured secondary diameter 43.8+/- 0.3 cm (17.24 +/- 0.12 inches); secondary obscuration 45.7+/- 0.3 cm (18+/-0.12 inches). Primary mirror is 41 inches diameter, center hole approximately 13.5 inches diameter, thickness is 6 inches

• Primary mirror support: there is a aluminum ”stovepipe” that is bolted to the back frame. It has anout diameter of 11.982-11.984 inches as measured by Rick from Sunspot. There is a blue plastic ring around this that is approximately 1 cm thick and 6 inches long. The mirror fits over this ring. The mirror rests on the ring on just the front and back faceplates, as the central part of the mirror is recessed. On the front plate, the hole is somewhat oversized when compared to the outer diameter of the plastic; measurements with shims gave the space between the mirror and the plastic as less than 0.050” everywhere, >0.001” from 1-9:00, >0.015” from 2-8:45, >0.020” from 2-8:00, >0.025” from 2-7:00, >0.030” from 2-6:30, and > 0.040” from 3-6:00, and had a maximum spacing of about 0.045” at 5:00, when measured in April 2004 (i.e. the mirror was shifted a bit to the right).

In early May 2004, we pulled the primary and removed the plastic sleeve. The sleeve was smoothed at the Sunspot machine lab, so the tolerances are now even greater than before. In addition, the sleeve was unable to go all the way to the bottom of the stovepipe because of the weld there, so the inner bottom part of the sleeve was ground away to allow the sleeve to go to the bottom, and hopefully, not get pushed around by thermal expansion at the weld.

• Rotator mounting details: 12(?) hole bolt pattern with 1/4-20 bolts spaced evenly on a circle 14 inches diameter. From the center of the rotator, you can mount anything out to a radius of 12 inches. Once you are back 6 inches from the rotator, there is some additional clearance; by the time you are 8 inches from the rotator, there are no constraints on a mounted instrument.

• Filter wheel mounting details.

• Alt-az-rotator system has 3 axes of motion; at current time, rotator exists on only one of the Nasmyth ports. Altitiude and azimuth axes have limit switches and magnetic home sensors. Rotator axis does not have limit switches, only a magnetic home sensor; to avoid rotator cable wrap problems, the rotator position must be carefully stored using software and remembered from session to session.

• All drives are friction drives. Altitude and azimuth have friction driven encoders as well; rotator only has the motor drive. On the azimuth axis, there is clear evidence that the “scale” of the axis, degrees/mm, varies slowly with temperature despite the fact that the material of the drive shaft and drive wheel are the same. We have plans to implement a temperature-dependent scale to correct for this to first order.

2 Instruments

2.1 Current guide box/filter wheel

We had a new guider and filter wheel built by Astronomical Consultants and Equipment (Peter Mack) out of Tucson. This was installed in mid-1999. The idea was to make a much thinner filter wheel which will allow the guider to be located much closer to the focal plane. This allows the guider to come closer on-axis without vignetting the primary beam to the science instrument. This is motivated so we can maximize the size of baffles in the system to minimize scattered light, which currently poses a problem.

The new guider consists of a guide box which is 9.25 inches thick, followed by a thin (1.313 inch) filter wheel, followed by the CCD. The guider module plugs into the side of the guider box, and uses our SpectraSource CCD. The guide module has a pickoff mirror to direct light to the guide camera. The entire pickoff/camera assembly is on a radial stage; in addition, the pickoff mirror is on a separate stage to allow focus control.

2.2 Old guide box

Old design had a spacer box mounted on rotator mount, which has a back mounting plate 5.25” behing the rotator mount. An off-axis guider camera is located inside this box, mounted to the back plate. This mount holds a 1.25” diagonal mirror along with a SpectraSource 512x512 CCD camera with 20 micron pixels. The center of the guider is currently located ∼ 2300 arcseconds off axis. However, at the current time, the diagonal mirror housing partially vignettes the beam going to the science instrument. There is currently no corrector in front of the guider.

Behind the spacer, a filter wheel is mounted. The current filter wheel is 3.5” thick. It accomodates 6 2” square filters.

Behind the filter wheel, the science CCD camera is located. It has a focal plane which is located 1.25 ” (optical distance) behind the mounting surface. The science CCD is 1024x1024 array put together by Princeton Instruments. The CCD pixels are 24 microns square.

The old design placed the CCD array about 2 inches closer to the secondary than is suggested by ray-tracing using the nominal optical parameters.

2.3 Apogee CCD camera

We are currently operating using an APOGEE AP7p camera as the science imager. This has a 512x512 thinned backside illuminated SITe chip with 24μ pixels in it. It is a thermoelectrically cooled camera which can operate about 50 degrees below ambient, leading to operating temperatures between -50 and -30 C, depending on the season. Dark current is significant at the warmer temperatures.

2.4 Princeton CCD camera

We have a Princeton Instruments CCD camera which has a 1024x1024 SITe array with 24μ pixels. This is a liquid nitrogen cooled system. We had a commandable fill system built for this dewar by VBS industries (see systems documentation).

The PI camera was purchased in the early 1990s. We have had a fair amount of trouble with low level noise in the camera, and it has been sent back to the manufacturer several times. Princeton Instruments has been taken over by Roper Scientific.

3 Computer and software control

The telescope, science CCD and filter wheel, and guider CCD are each controlled by a separate PC. The telescope is controlled by the TOCC computer, running DOS, the science CCD by the PI computer running Windows 3.11, and the guider CCD by the SPEC computer running DOS.

An original version of the alt-az control system running on the TOCC computer was written by the AutoScope corporation, but this had many problems and large portions of it have been rewritten (by Jon Holtzman) to form the current version. This software also controls the dome and dome shutter, as well as commandable fill system for the CCD dewar.

The science CCD is controlled by software provided by Princeton Instruments.

The guide CCD is controlled by software written by Jon Holtzman using library routines provided by SpectraSource.

All three programs look, in addition to input from the local keyboard(s), for input in specified files which are visible over the network. This allows for complete remote operation.

For remote operation, a suite of programs (tcomm) is run from a UNIX workstation. Basically, a program is run which talks with each of the three computers in the dome. In addition, there is a master command program which is the user interface and sends commands to the UNIX client programs. There is also a status program which monitors status files sent back from the telescope and the CCD programs and displays a status window with the current pointing of the telescope, exposure/filter information, etc. Finally there is a program which controls the power to various devices in the dome via a network power switching device.

3.1 Notes on pointing software

We define the rotator coordinate system to be a right-handed coordinate system centered on the rotator center, with axes which will be defined to run parallel to NS and EW when the rotator position angle is set to 0. Call rotator coordinates (x′, y′). Right ascension and declination (sky coordinates) are related by the fact that the sky coordinates are left-handed and by the rotator position angle (ω):

−α = x′ cos ω − y′ sin ω
δ = x′ sin ω + y′ cos ω

where the sign convention for the position angle is that a positive position angle rotates N counter clockwise on the detector when viewed in sky orientation.

Any instrument will then be defined by its pixel scale (S_{x}, S_{y}), rotation of the pixels relative to the rotator coordinate system (θ), and the location of the detector center in the rotator coordinate system (c_{x}, c_{y}). If instrument coordinates are represented by (x, y), then

x′ = (x − x_{c})S_{x} cos θ + (y − y_{c})S_{y} sin θ + c_{x}
y′ = (x − x_{c})S_{x} sin θ − (y − y_{c})S_{y} cos θ + c_{y}

S and θ are determined by moving a star around in the detector (command INIT, GINIT). In general, S_{x} and S_{y} will be constrained to have the same absolute value (or to have a ratio of absolute values given by the ratio of pixel dimensions); if the detector has a left-handed orientation in sky coordinates (i.e. origin at upper left for sky orientation), the scales will have the same sign, but if right-handed, the convention is to make S_{x} negative.

c_{x} and c_{y} are related to the offset which is reqired to move a star from the center of the rotator to the center of the instrument; if a QM A B moves the star from the center of the rotator to the center of the instrument, then c_{x} = A and c_{y} = −B, where there is a sign flip for right ascension since sky coordinates are left-handed while rotator coordinates are right-handed.

Pointing models are determined to put stars at the rotator center.

To place an object in the center of a desired instrument, one thus needs to add an instrument correction to the (α, δ) of the target. Since we need to move in the opposite direction of the offset, we add:

Δα = c_{x} cos ω − c_{y} sin ω
Δδ = −c_{x} sin ω − c_{y} cos ω

To offset the telescope in instrument coordinates, one has

Δx′ = ΔxS_{x} cos θ + ΔyS_{y} sin θ
Δy′ = ΔxS_{x} sin θ − ΔyS_{y} cos θ

or

Δα = −ΔxS_{x} cos(θ + ω) − ΔyS_{y} sin(θ + ω)
Δδ = +ΔxS_{x} sin(θ + ω) − ΔyS_{y} cos(θ + ω)

Remember, if one wishes to move an object a desired number of pixels, one must move the telescope in the opposite direction.

3.2 Notes on guider software

The guider has two parallel stages, a positioning stage and a focus stage. The positioning stage moves both the camera and the pickoff mirror, while the focus stage moves the pickoff mirror only. Hence, changing the focus also changes the position in the sky.

Let P be the radial position in the sky in arcsec, x be the radial position in microns, R the R-stage step position, and F the F-stage step position. Let dR/dF be the ratio of R-steps to F-steps which keeps P constant. Let dP/dR be the scale of R-stage, i.e. arcsec/steps; dP/dR = Sdx/dR, where S is the telescope focal plane scale. Then

P = [(R − R_{0}) + (F − F_{0})\frac{dR}{dF}]\frac{dP}{dR} + P_{0}

However, this is complicated by the fact that the guider focus changes with position in the sky because the focal plane is curved. If we parameterize this change by:

F (P) = F_{0} + k_{1}P + k_{2}P^{2}

then

P = [(R − R_{0}) + (k_{1}P + k_{2}P^{2})\frac{dR}{dF}]\frac{dP}{dR} + P_{0}

Inverting to get R(P),

(R − R_{0})\frac{dP}{dR} = (P − P_{0}) − (k_{1}P + k_{2}P^{2})\frac{dR}{dF} \frac{dP}{dR}

or

R = \frac{(P − P_{0}) − (k_{1}P + k_{2}P^{2})\frac{dR}{dF} \frac{dP}{dR}}{\frac{dP}{dR}} + R_{0}

4 Telescope calibration items

After changing the optical configuration of the telescope, several calibration steps may need to be taken:

1. Determine rotator center. Probably best done by going to star field, set position angle, disable rotator (ZDISABLE), change position angle by less than 180 degrees, avoiding cable wrap issues, do a long (300-600s) integration and find the center.

2. Do a pointing model. Put rotator center in pm.pro and do model using recentering. This creates pm???.fits files. Procedure posrecen.pro will turn these into TPOINT input file. TPOINT generates a pointing model: INDAT, CALL ALTAZ, USE NRX NRY, FIT, OUTMOD.

3. Do instrument blocks. Need to determine offset from rotator center to center of each instrument at PA=0 with QM commands. Guider CX coordinate currently needs to be compiled into ccd/gccd.c. GINIT should determine scales and rotations, which together with offsets form instrument blocks.

4. Determine proper guider focus (at rotator==0?) and set guider home position accordingly.

5. Focus offsets, with altitude and rotator position/stage location for guider: foctesta.pro, foctestg.pro, focadj.pro. Focus offsets for filters.